Please help...
If a pro basketball player has a vertical leap of about 20 inches, what is his hang time? Use the hang time function V=48T^2

Answer by richwmiller(17219)

plug in 20 for v and solve for t
20=48t^2
20/48=t^2
5/12=t^2
sqrt(5/12)=t